OpenStudy (rahulmr):

Will fan and give medal. The perimeter of a circular sector is 20 cm and the radius is increasing at 5 cm/s. At what rate is the sector changing when the radius is 10 cm. Also at what rate is the area changing when the radius is 10 cm. Hi, Can anyone help me in this question. I don't know where should i start. Thanks.

OpenStudy (devonhoward15):

nice car

OpenStudy (isaiah.feynman):

Hmmm...

OpenStudy (isaiah.feynman):

The area of a sector is given as \[A = \frac{ r^{2} \theta }{2}\]

OpenStudy (isaiah.feynman):

To find the rate at which the area is increasing, you will have to implicitly differentiate the formula for area. Note that theta is in radians.

OpenStudy (rahulmr):

Thanks for that @Isaiah.Feynman , but don't you think i should use \[P=2*R+R*\theta \]

OpenStudy (isaiah.feynman):

I don't see how. I don't quite understand what is meant by "At what rate is the sector changing when the radius is 10cm"

OpenStudy (isaiah.feynman):

Also the perimeter would be \[P = (2+\theta)r\]

OpenStudy (rahulmr):

What i did was \[p=2r+r \theta \] since p=20cm \[r=20/(2+\theta) \] then i differentiate the equation and i got \[-20/(2+\theta)^2 \]

OpenStudy (rahulmr):

now i don't know what should i do next

OpenStudy (isaiah.feynman):

I think you should differentiate first before plugging in numbers.

OpenStudy (freckles):

@Isaiah.Feynman is right at what rate is the sector changing is very vague what specifically are they talking about about the sector